OpenStudy (anonymous):

solve 3x^2 +4X^2 I keep getting 25 but that isn't the right answer

OpenStudy (x3_drummerchick):

is it a factoring problem?

OpenStudy (anonymous):

\[2x^2+4x^2\] is not an equation to solve you can add and get \[2x^2+4x^2=6x^2\]

OpenStudy (triciaal):

typo should be 3x^2 + 4x^2 = 7x^2

OpenStudy (triciaal):

@nikita17 are you sure you copied all of the problem? usually to solve you are given an equation.

OpenStudy (anonymous):

Really! I was doing 3x^2 as 3*3 =9 and 4x^2 as 4*4=16 which combined would equal 25 even the calculator agreed but it was wrong

OpenStudy (anonymous):

This was just combining like terms these were the only 2 like terms

OpenStudy (anonymous):

\[3x^2=3\] is not always true you have a variable there \[3x^2\] is three times some number \(x\) squared

OpenStudy (anonymous):

was there a question that came along with \[3x^2+4x^2\]?

OpenStudy (anonymous):

yes combine like terms

OpenStudy (anonymous):

ok now we have a question and now we can answer it but before we do, what is three hundred plus four hundred?

OpenStudy (anonymous):

I think triciaal was correct she said 7x^2

OpenStudy (anonymous):

700

OpenStudy (anonymous):

yes and what is three hamburgers and four hamburgers?

OpenStudy (anonymous):

or \[3\heartsuit +4\heartsuit\]?

OpenStudy (triciaal):

@satellite73 was correct he just had the wrong number he used 2 instead of 3

OpenStudy (anonymous):

yes i see what you are getting at satellite 73

OpenStudy (anonymous):

Yes I see now

OpenStudy (anonymous):

so it is clear right? \[3x^2+4x^2=7x^2\] on account of \(3+4=7\)

OpenStudy (anonymous):

It was confusing because when I put the equation in my calculator it was 25 and I could also see how it came up with that

OpenStudy (anonymous):

my dear this is not a calculator exercise it has nothing to do with your calculator

OpenStudy (anonymous):

Yes thank you both satellite73 and triciaal

OpenStudy (triciaal):

remember the calculator only does what you tell it to do

OpenStudy (anonymous):

as a matter of fact, i am willing to bet that you cannot "put the equation in the calculator"

OpenStudy (anonymous):

our teacher had us download a special type of calculator

OpenStudy (anonymous):

I believe triciaal is right. the calc will do what you tell it to do

OpenStudy (anonymous):

Thanks again both of you =)

OpenStudy (triciaal):

you are welcome